# Singular equivalence and singular locus

### v1

@@ Abstract.
For a commutative Gorenstein Noetherian ring $R$, we construct an affine scheme $X$ solely from singularity category $D_{sg}(R)$ of $R$ such that there is a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$. As an application, for two such rings with equivalent singularity categories, we prove that the singular loci in their affine schemes have the same dimension.
@@ 1.2. Main result
In [20], Orlov proved that the completion of a variety along its singular locus determines its singularity category, up to the idempotent completion of a triangulated category. We want to answer the inverse question of Orlov’s result up to some extent. Unfortunately, we need some restrictive conditions on the commutative ring. Our main result is the following.
@@
Let $R$ be a commutative Gorenstein Noetherian ring. Then there is an affine scheme $X$ constructed solely from the triangulated category $D_{sg}(R)$, and a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$.
@@ 1.3. Idea of the proof
    At last, with some preparations (see Lemmas 4.1 and 4.2), we construct scheme $X$ and morphism $X\rightarrow\mathrm{Spec}(R/I)$, and then prove Theorem 1.1.
The paper is organized as follows. In Section 2, we introduce the notions and some properties of singularity categories, singular equivalences and DG categories. In Section 3, we prove the statements in the steps (1) and (2) above. In Section 4, we give the proof of the statement in step (3), and get the main result.
@@ 2.1. Singularity category and singular equivalence
There are many examples of singular equivalence. For example, Chen and Sun introduced the singular equivalence of Morita type in [8]. More generally, Wang introduced singular equivalence of Morita type with level in [23]. We give some examples of singular equivalence.
@@
mapping $v$ to zero. Notice that $S[u]/(f)$ is an $\big{(}S[u,v]/(f+uv)\big{)}\otimes S^{op}$-module.
@@
It follows that $S[u]/(f)\in\mathrm{Perf}\big{(}S[u,v]/(f+uv)\big{)}$. Thus, there is a triangulated functor
@@
In [16], Knorrer proved that the above triangulated functor is a singular equivalence between $S/(f)$ and $S[u,v]/(f+uv)$.
Example 2.4.
Let $R$ and $S$ be two commutative Noetherian rings, where $R$ is homologically smooth. Let $N$ be an $S\otimes R^{op}$-module such that it is projective over $S$. Set
$A:=\left(\begin{matrix}R&N\\ 0&S\end{matrix}\right),$
and $M:=\left(\begin{matrix}N\\ S\end{matrix}\right)$. Then $M$ is an $S\otimes A^{op}$-module and a projective $S$-module. In [7, Theorem 4.1], Chen showed that the triangulated functor
$D_{sg}(A)\xrightarrow{(-)\otimes^{\mathbb{L}}_{A}M}D_{sg}(S).$
is a singular equivalence. Furthermore, $M$ is also a projective $A^{op}$-module. The triangulated functor
$D_{sg}(S)\xrightarrow{(-)\otimes^{\mathbb{L}}_{S}M^{T}}D_{sg}(A)$
is a singular equivalence, where $M^{T}$ is the transpose of $M$.
2.2. DG category and its DG enhancement
@@
    $\mathrm{Hom}_{\mathrm{H}(\mathcal{D})}(X,Y)=\mathrm{H}^{\ast}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}=:\bigoplus\limits_{i\in\mathbb{Z}}\mathrm{H}^{i}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}$, for any objects $X,Y$.
@@ A DG category $\mathcal{C}$ is called pretriangulated if the essential image of the functor
Definition 2.8.
Let $\mathrm{D}$ be a triangulated category.
(1) A DG enhancement of $\mathrm{D}$ is a pair $(\mathcal{D},\psi)$, where $\mathcal{D}$ is a pre-triangulated DG category and $\psi:\mathrm{H}^{0}(\mathcal{D})\rightarrow\mathrm{D}$ is an exact equivalence.
(2) $\mathrm{D}$ has a unique DG enhancement if for two DG enhancements $(\mathcal{D},\psi)$ and $(\mathcal{D}^{\prime},\psi^{\prime})$, there exists a quasi-functor $F:\mathcal{D}\rightarrow\mathcal{D}^{\prime}$ such that $\mathrm{H}^{0}(F)$ is an equivalence.
In the definition, the equivalence of $\mathrm{H}^{0}(F)$ means that $F$ is an equivalence in $\mathrm{Hqe}$.
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
    $\mathrm{H}(\pi)$ gives a triangulated functor $\tilde{\mathcal{A}}^{tr}\rightarrow\mathcal{C}^{tr}$, which induces an equivalence:
@@
From [9, §3], we know that, for any DG category $\mathcal{A}$ and its full DG subcategory $\mathcal{B}\subset\mathcal{A}$, the DG quotient of $\mathcal{A}$ by $\mathcal{B}$ exists. On the other hand, it is well known that both $D^{b}(A)$ and $\mathrm{Perf}(A)$ admit a canonical DG category structure. Here, we denote by $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ the DG categories corresponding to $\mathrm{Perf}(A)$ and $D^{b}(A)$ respectively. Moreover, both $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ have canonical pre-triangulated structures. Hence, by taking the DG quotient, $D_{sg}(A)$ is endowed with a DG category structure which is a DG enhancement for $D_{sg}(A)$. We denote by $S_{dg}(A)$ this DG category associated to $D_{sg}(A)$.
In [15, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
@@
Next we recall the following proposition (see [9, 1.3.1]).
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [14, Theorem 8.1], the triangulated functor
@@
By the definition of tensor product of DG modules over a DG category (see [17, 3.5]), we know that
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [17, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
@@
3.2. Hochschild cohomology of singularity category
@@ 4. Proof of the main result
In this section, we prove Theorem 1.1. First, we prove the following.
Lemma 4.1.
Let $R$ be a commutative Gorenstein Noetherian ring. Then $D_{sg}(R)$ has a unique DG enhancement.
Proof.
In [22], Rouquier showed that, for any separated scheme $Z$ of finite type over $k$, $D^{b}(Z)$ admits a strong generator. Meanwhile, we know that any strong generator must be a classical generator. Thus suppose $Z$ is an affine scheme $\mathrm{Spec}(R)$, then $D^{b}(R)$ admits a classical generator, denoted by $E$.
Next, the natural localization functor $\pi:D^{b}(R)\rightarrow D^{b}(R)\big{/}\mathrm{Perf}(R)=D_{sg}(R)$ maps $E$ to a classical generator $\pi(E)$ of $D_{sg}(R)$. Meanwhile, since $R$ is Gorenstein, we have
$D_{sg}(R)\simeq\underline{\mathrm{CM}}(R)$
as triangulated categories. Then $D_{sg}(R)$ is an algebraic triangulated category. Thus we have the following triangle equivalence:
$D_{sg}(R)\xrightarrow{\sim}\mathrm{Perf}(T),$
where $T:=\mathbb{R}\mathrm{Hom}_{D_{sg}(R)}(E,E)$ (see [14]).
Note that $T$ is a differential graded algebra. By applying [18, Theorem 2.8], we know that $D(T)^{c}=\mathrm{Perf}(T)$ has a unique DG enhancement, where $(D(T))^{c}$ is the full subcategory, consisting of compact objects, of $D(T)$. It implies that $D_{sg}(R)$ has a unique DG enhancement which is given by the Drinfeld quotient. We completed the proof. ∎
By the above lemma, we know that $S_{dg}(R)$ is the unique DG enhancement of $D_{sg}(R)$.
Lemma 4.2.
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a singular equivalence
$F:D_{sg}(R_{1})\rightarrow D_{sg}(R_{2}).$
Then there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\rightarrow D_{sg}(R_{2}),$
where $\tilde{F}$ is a lift of $F$, i.e., $F=\mathrm{H}(\tilde{F})$.
Proof.
By Lemma 4.1, $\mathrm{Perf}(T)$ has a unique DG enhancement. Thus, by Definition 2.8, there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\rightarrow S_{dg}(R_{2})$
such that it is a lift of $F$, i.e., $F=\mathrm{H}(\tilde{F})$ (see [2, 1.15]). QED. ∎
Now we are ready to show the following.
@@
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a singular equivalence
$D_{sg}(R_{1})\rightarrow D_{sg}(R_{2}).$
@@
By Lemma 4.2, we obtain that there is a quasi-equivalence
$\tilde{F}:S_{dg}(R_{1})\xrightarrow{(-)\otimes^{\mathbb{L}}_{R_{1}}M}S_{dg}(R_{2})$
induced by the triangulated functor $F$. Thus by Theorem 3.3, their Hochschild cohomologies are isomorphic. It implies that
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions.

### v2

# DG Singular equivalence and singular locus

@@ Abstract.
For a commutative Gorenstein Noetherian ring $R$, we construct an affine scheme $X$ solely from DG singularity category $S_{dg}(R)$ of $R$ such that there is a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$. As an application, for two such rings with equivalent DG singularity categories, we prove that the singular loci in their affine schemes have the same dimension.
@@ 1.2. Main result
In [21], Orlov proved that the completion of a variety along its singular locus determines its singularity category, up to the idempotent completion of a triangulated category. We want to answer the inverse question of Orlov’s result up to some extent.
Unfortunately, for general commutative Gorenstein Noetherian ring $R$, the information of the triangulated category $D_{sg}(R)$ is not enough to detect the geometry that we would expect. We have to consider some DG enhancement of $D_{sg}(R)$. The singularity category $D_{sg}(R)$ admits a canonical DG enhancement given by the DG quotient $S_{dg}(R):=\mathcal{D}^{b}(R)/\mathcal{P}\mathrm{erf}(R)$ ([10, §3]), where $\mathcal{D}^{b}(R)$ is the canonical DG enhancement of $D^{b}(R)$, which then induces a DG enhancement $\mathcal{P}\mathrm{erf}(R)$ of $\mathrm{Perf}(R)$. Our main result is the following.
@@
Let $R$ be a commutative Gorenstein Noetherian ring. Let $S_{dg}(R)$ be the triangulated category described as above. Then there is an affine scheme $X$ constructed solely from $S_{dg}(R)$, and a finite surjective morphism $X\rightarrow\mathrm{Spec}(R/I)$, where $\mathrm{Spec}(R/I)$ is the singular locus in $\mathrm{Spec}(R)$.
@@ 1.3. Idea of the proof
    At last, we construct scheme $X$ and morphism $X\rightarrow\mathrm{Spec}(R/I)$, and then prove Theorem 1.1.
The paper is organized as follows. In Section 2, we introduce the notions and some properties of DG categories, DG singularity categories and DG singular equivalences. In Section 3, we prove the statements in the steps (1) and (2) above. In Section 4, we give the proof of the statement in step (3), and get the main result.
@@ 2.1. Singularity category and singular equivalence
There are many examples of singular equivalence. For example, Chen and Sun introduced the singular equivalence of Morita type in [8]. More generally, Wang introduced singular equivalence of Morita type with level in [24].
@@
mapping $v$ to zero. Notice that $S[u]/(f)$ is an $\big{(}S[u,v]/(f+uv)\big{)}\otimes(S/(f))^{op}$-module.
@@
It follows that $S[u]/(f)\in\mathrm{Perf}\big{(}S[u,v]/(f+uv)\big{)}$. Thus, there is a triangle functor
@@
In [17], Knorrer proved that the above triangle functor is a singular equivalence between $S/(f)$ and $S[u,v]/(f+uv)$.
2.2. DG category and DG functor
@@
    $\mathrm{Hom}_{\mathrm{H}(\mathcal{D})}(X,Y)=\mathrm{H}^{\ast}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}:=\bigoplus\limits_{i\in\mathbb{Z}}\mathrm{H}^{i}\big{(}\mathrm{Hom}_{\mathcal{D}}(X,Y)\big{)}$, for any objects $X,Y$.
@@ 3.1. Generalized Tate-Hochschild complex of associative algebra
    $\mathrm{H}(\pi)$ gives a triangle functor $\tilde{\mathcal{A}}^{tr}\rightarrow\mathcal{C}^{tr}$, which induces an equivalence:
@@
From [10, §3], we know that, for any DG category $\mathcal{A}$ and its full DG subcategory $\mathcal{B}\subset\mathcal{A}$, the DG quotient of $\mathcal{A}$ by $\mathcal{B}$ exists. On the other hand, it is well-known that both $D^{b}(A)$ and $\mathrm{Perf}(A)$ admit a canonical DG category structure. Here, we denote by $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ the DG categories corresponding to $\mathrm{Perf}(A)$ and $D^{b}(A)$ respectively. Moreover, both $\mathcal{P}\mathrm{erf}(A)$ and $\mathcal{D}^{b}(A)$ have canonical pre-triangulated structures. Hence, by taking the DG quotient, $D_{sg}(A)$ is endowed with a DG category structure which is a DG enhancement for $D_{sg}(A)$. We denote by $S_{dg}(A)$ this DG category associated to $D_{sg}(A)$.
Definition 3.3.
Let $A$ and $B$ be two associative algebras. $A$ and $B$ are called DG singular equivalent if there is a quasi-equivalence
$S_{dg}(A)\xrightarrow{\sim}S_{dg}(B).$
It is well-known that the DG singular equivalence between $S_{dg}(A)$ and $S_{dg}(B)$ implies that $D_{sg}(A)\simeq D_{sg}(B)$. But, in general, singular equivalence cannot be lifted to DG singular equivalence. However, we have the following proposition for some spacial case.
Proposition 3.4 ([9, Proposition 3.1]).
Keep the settings as above. Let $M$ be a $B\otimes A^{op}$-module which is a projective $B$-module and projective $A^{op}$-module. Then the following statements are equivalent.
  1. (1)
    The triangle functor
    $(-)\otimes^{\mathbb{L}}_{A}M:D_{sg}(A)\rightarrow D_{sg}(B)$
    is an equivalence;
  2. (2)
    The quasi-functor
    $(-)\otimes^{\mathbb{L}}_{A}M:S_{dg}(A)\rightarrow S_{dg}(B)$
    is a quasi-equivalence.
Here, we call singularity equivalence given by some $B\otimes A^{op}$-module $M$ as above proposition singularity equivalence of Morita type ([8]).
Example 3.5.
We continue the study of Example 2.3. In the argument of Example 2.3, we know that $M$ is a projective $S[u,v]/(f+uv)$-module, where $M:=S[u]/(f)$. Moreover, it is easy to check that $M$ is a projective $S^{op}$-module. Hence, we get a triangle functor
$(-)\otimes^{\mathbb{L}}_{A}M:D_{sg}(S)\rightarrow D_{sg}\big{(}S[u,v]/(f+uv)\big{)}.$
By Knorrer’s periodicity theorem, this functor is a triangle equivalence. Finally, by Proposition 3.4, we get that it is in fact an equivalence
$S_{dg}(S)\simeq S_{dg}\big{(}S[u,v]/(f+uv)\big{)}.$
In [16, Theorem 1.1], Keller realized the Tate-Hochschild cohomologies of an algebra $A$ as the Hochschild cohomologies of $S_{dg}(A)$ in the following theorem.
@@
Next we recall the following proposition (see [10, 1.3.1]).
@@
sending the unique object of $A^{op}$ to $A$-module complex $A$. By the result of Keller in [15, Theorem 8.1], the triangle functor
@@
By the definition of tensor product of DG modules over a DG category (see [18, 3.5]), we know that
@@
If we replace $N^{\ast}$ by its projective $A$-module resolution, say, $P_{N^{*}}$, then by [18, 3.5], $h_{N^{\ast}}\otimes_{\mathcal{P}\mathrm{erf}(A)}^{\mathbb{L}}\tilde{h}_{M^{\ast}}$ is isomorphic to $h_{P_{N^{\ast}}}\otimes_{\mathcal{P}\mathrm{erf}(A)}\tilde{h}_{M^{\ast}}$. It follows that
@@
3.2. Hochschild cohomology of DG singularity category
@@ 4. Proof of the main result
In this section, we prove Theorem 1.1.
@@
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a DG singular equivalence
$F:S_{dg}(R_{1})\xrightarrow{\sim}S_{dg}(R_{2}).$
@@
In the meantime, there is a quasi-equivalence
$F:S_{dg}(R_{1})\xrightarrow{\sim}S_{dg}(R_{2}).$
Thus by Theorem 3.6, their Hochschild cohomologies are isomorphic. It implies that
@@
By Proposition 3.4, any singularity equivalence of Morita type gives a DG singularity equivalence, and thus we get the following.
Corollary 4.2.
Let $R_{1}$ and $R_{2}$ be two commutative Gorenstein Noetherian rings. Suppose that there is a singular equivalence of Morita type
$\Phi:D_{sg}(R_{1})\xrightarrow{(-)\otimes^{\mathbb{L}}_{R_{1}}M}D_{sg}(R_{2}),$
for some $R_{2}\otimes R^{op}_{1}$-module $M$. Then the two singular loci in the affine schemes of these two rings have the same dimension.
@@ Acknowledgements.
We would like to thank Xiaojun Chen and Huijun Fan for their encouragement, support and suggestions. Moreover, we appreciate Martin Kalck for pointing out an error in the previous version of this paper.